OpenStudy (anonymous):

Solve this equation 17 = –13 – 8x

OpenStudy (anonymous):

30/8=15/4=3.75

OpenStudy (vishweshshrimali5):

30 = 8x => x = 30/8 = 15/4

OpenStudy (anonymous):

decimals